OpenStudy (anonymous):

Evaluate the expression: 6^7/6^3

OpenStudy (anonymous):

Does it equal 1296

OpenStudy (anonymous):

Yes, when diving two exponents with the same base you substract the exponent: \[{a^m \over a^n} = a^{m - n}\] 6^7/6^3 = 6^4 = 1296

OpenStudy (anonymous):

or is it 6^21

OpenStudy (anonymous):

You're dividing two exponents, \[{6^7 \over 6^3} = \frac{6*6*6*6*6*6*6}{6*6*6}\] You can cancel three of them out \[\frac{6*6*6*6\cancel{*6*6*6}}{\cancel{6*6*6}} = 6*6*6*6 = 6^4 = 1296\] And in general, for a number a, \[\Large \frac{a^m}{a^n} = a^{m-n}\]

OpenStudy (anonymous):

Oh, okay. Thank you :)